OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (4, −7) and a directrix of y = −15. Put the equation in standard form

OpenStudy (anonymous):

f(x) = one sixteenthx2 − 8x + 11 f(x) = one sixteenthx2 − 8x − 10 f(x) = one sixteenthx2 − one halfx + 11 f(x) = one sixteenthx2 − one halfx − 10

OpenStudy (campbell_st):

well the distance between the focus and directrix is twice to focal length so from y = -7 to y = -15 is 8 units... so the focal length is 4 the vertex is a focal length below the focus...  so know you know the focal length and vertex, you can find the equation, hope this helps

OpenStudy (campbell_st):

then the equation is \[y = \frac{1}{4a}(x - h)^2 + k\] you know the value of a, the focal length and you know the vertex (h, k) just substitute and evalute